Two forces act on a 2-kg box. The first force is 10 N to the right, while the second force if 5 N to left. Determine the resultant acceleration.

1
Expert's answer
2021-11-07T17:07:37-0500

The resultant force is


"F_\text{net}=F_1-F_2=10-5=5\text{ N}."

The acceleration:


"a=F_\text{net}/m=10/2=2.5\text{ m/s}^2."